OpenStudy (anonymous):

Can someone check? Please and thank you! An ellipse has foci at (–1, 1) and (–1, –7). The major axis is 12 units long. Show work for all answers. -Determine the center of the ellipse. -Determine the vertices of the ellipse. -Find the equation of the ellipse. I got (-1, -3) as the center. (6,0) and (-6, 0) for the vertices and y^2/ 36 + X^2/ 20 = 1

mathslover (mathslover):

Sorry not so good in this topic :'(

OpenStudy (anonymous):

its okay! thanks anyways! :)

OpenStudy (anonymous):

and i just did these types of problems in summer school... smh

OpenStudy (anonymous):

haha am i doing them right?

OpenStudy (anonymous):

so the major axis is y, 6 high and 6 low...thinking

OpenStudy (anonymous):

i m unsure about the vertices!

OpenStudy (anonymous):

hold on its coming back to me lol

OpenStudy (anonymous):

hahaha!

OpenStudy (anonymous):

this graph is shifted to the left -1 so shouldnt your answer be in the (y+1)^2 + (x+1)^2

OpenStudy (anonymous):

form

OpenStudy (anonymous):

i always get confused about that! but i think it said its in x^2/ a^2 form? which one makes sense?

OpenStudy (anonymous):

if your center was at (0,0) then it would be x^2 + y^2 = 1, but this is shifted to it has to be the other form

OpenStudy (anonymous):

oh ok! so how can i change it to the other form?

OpenStudy (anonymous):

simply change the numerator to (x+1)^2

OpenStudy (anonymous):

you might leave the y^2 alone i cant remember

OpenStudy (anonymous):

id literally have to dig through some of my notes to get you the right answer tho lol

OpenStudy (anonymous):

oh haha i wish i had some notes i could look at!

OpenStudy (anonymous):

i might be able to send you something hold up

OpenStudy (anonymous):

may I share Salman Khan's video tutorial on focuses of ellipse?

OpenStudy (anonymous):

loloololololol thanks guys!

OpenStudy (anonymous):

http://www.youtube.com/watch?v=QR2vxfwiHAU

OpenStudy (anonymous):

center will be (-1,-3) vertices (-1,3) (-1,-9)

OpenStudy (anonymous):

nick, thanks! but no thanks

OpenStudy (anonymous):

what does this have to do with you? lol

OpenStudy (anonymous):

it is not my homework to solve :)

OpenStudy (anonymous):

thank you so muchhh! ive been trying to work it out forever!

OpenStudy (anonymous):

i dont think thew equation is right! cause it says b = 20

OpenStudy (anonymous):

?

OpenStudy (anonymous):

cause (X+1)^2/ 20

OpenStudy (anonymous):

im not sure yet of how to make the equation out of that. do u know?

OpenStudy (anonymous):

well the equation is (y+3)^2/ 36 + (X+1)^2/ 20 = 1 im just not sure that 20 is ther right number

OpenStudy (anonymous):

so this (y+3)^2/ 36 + (X+1)^2/ 20 = 1 is the confirmed answer?

OpenStudy (anonymous):

yes but im not sure about the 20! well i think its the confirmed answer haha

OpenStudy (anonymous):

well a = 6, c=4, then u use b^2=a^2-c^2

OpenStudy (anonymous):

your left with b^2=20

OpenStudy (anonymous):

its c^2 = a^2- b^2

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

no your right

OpenStudy (anonymous):

yesah but then it doesnt work!

OpenStudy (anonymous):

it does work

OpenStudy (anonymous):

add the b^2 subtract the c^2

OpenStudy (anonymous):

when its plugged in you have 16=36-b^2

OpenStudy (anonymous):

16- 36

OpenStudy (anonymous):

add the b^2

OpenStudy (anonymous):

-20

OpenStudy (anonymous):

even then you have -20=-b^2, you would have to divide both sides by negative 1 to make the b positive

OpenStudy (anonymous):

you gotta believe! lol

OpenStudy (anonymous):

yeah so it would still be 20! haah i will thanks! I OWE YOU! YOU HONESTLY SAVED MY LIFE!

OpenStudy (anonymous):

good i can sleep peaceful now lol good luck

OpenStudy (anonymous):

HAHA SORRY !!